Find the smallest number θ larger than 11π such that cos θ = -1/2.
Give exact answer

Answer by lwsshak3(11628)

Find the smallest number θ larger than 11π such that cos θ = -1/2.
Give exact answer.
***
arccos(-1/2)=2π/3+2πn,n=integer
2π/3+2πn>11π
2+6n>33
6n>31
n>31/6=5+
n=6
..

smallest number θ larger than 11π = 38π/3